A toy rocket is launched from the ground with an initial velocity of 80 feet per second. The height h of the rocket after t seconds is given by the equation h=80t-16t^2. How long after it is launched will the rocket be at its highest point?
Answer by stanbon(75887)

A toy rocket is launched from the ground with an initial velocity of 80 feet per second. The height h of the rocket after t seconds is given by the equation h=80t-16t^2. How long after it is launched will the rocket be at its highest point?
--------------
That equation is a quadratic with a=-16,b=80
Max occurs when t = -b/2a = -(80)/(-32) = 5/2 seconds
=================
Cheers,
Stan H.